Please help me solve this prolem. I just don't get it if there are not numbers.
Solve for P: A=P+Pr

Answer by checkley77(12844)

Letters should be easier because you don't have to use much math to solve the equations.
A=P+Pr
a=p(1+r)
p=a/(1+r)